You are here: Foswiki>Commissioning Web>Com20100706 (06 Jul 2010, PascaleHibon)Edit Attach

Commissioning: 6 July 2010 UT

Observer: Pascale H (LBTO) SW Group, Doug M (Tucson)
Telescope Operator: Steve A (LBTO)
Telescope Support: (LBTO)
Instrument Support: John M , Doug O (LBTO)
Adaptive Secondary Operator: Pascale H (LBTO), Armando R (Firenze)

Summary

Night was very clear with an average seeing of ~0.75-0.8"
Pointing model data.
Collimation data
Collection of data for Astigmatism problem.
Transformation data.
Dice5 data.

Preparations

Single frame names refer to: irtc.20100705.xxxxxx.fits

Cube names refer to: irtc.20100705_00xxx_cube.fits

Plan for the night: follow Doug's verification plan, in particular to transformation data, collimation and pointing models.

Details

Open

02:23 Open T_air=12C, Humidity=32%, Wind=6.5m/s
02:25 The shell is set as flat

Preset + Collimation

03:00 pointto ACT0343 Tel="ACQUIRE" gs=0 PA=0 PAM- Elevation 65deg
  • Astigmatism ~3arcsec
CA=6 IE=-5
03:10 pointto BS9162 Tel="ACTIVE" gs=0 Elevation 69deg
Steve has to restart GCS
03:13 pointto BS9162 Tel="ACTIVE" gs=0
  • Astigmatism ~3arcsec
Seeing ~ 0.8"

03:20 pointto BS9131 Tel="ACTIVE" gs=0 Elevation 29deg
  • wfscimage000010.fits - wfscimage000025.fits

Pointing model data
03:34 PCS turned on
  • location : /home/telescope/20100706.ptlog
03:35 gopoint exp-time=0.02s Nexp=500 delay=100s ploop+ log=gopointjuly062010
  • ACT0356 : Elevation 69deg Astigmatism ~8arcsec Saturated on WFS
03:39 The integration time of the WFS has been changed to 5000ms.
03:41 gopoint exp-time=0.02s Nexp=500 delay=100s ploop+ log=gopointjuly062010
  • 034414 cube214 : FAILED due ton strong Astigmatism. : ~6arcsec - Elevation 69
  • 034912 cube215 : FAILED due to star not found - Elevation 22deg
  • 035327 cube216 : SATURATED on IRTC Elevation 38deg
03:55 gopoint exp-time=0.005s Nexp=2000 delay=100s ploop+ log=gopointjuly062010
  • 035734 cube217 : FAILED Astigmatism~4arcsec - Elevation 63deg
04:01 gopoint exp-time=0.02s Nexp=500 delay=100s ploop+ log=gopointjuly062010

Pointing + Astigmatism test

04:14 pointto M5_0531 Tel="TRACK" gs=0 PA=0 Elevation 64deg
04:15 takeseq 0.01 30
  • 041838 cube 218
04:19 pointto M5_0699 Tel="TRACK" gs=0 PA=0 Elevation 57deg
04:22 takeseq 0.001 30
  • 042250 cube 219
Remark : the pointing is bad. 04:24 pointto M5_0912 Tel="TRACK" gs=0 PA=0 Elevation 40deg
04:26 pointto M5_0912 Tel="ACQUIRE" gs=0 PA=0
Steve is doing ptreset - ptspiral
04:33 pointto M5_0912 Tel="ACQUIRE" gs=0 PA=0 PAM-
CA=-2 IE=-24
04:36 pointto M5_0912 Tel="TRACK" gs=0 PA=0
04:37 takeseq 0.005 30
  • 041838 cube 220
04:39 pointto M5_0425 Tel="TRACK" gs=0 PA=0 Elevation 29deg
Star not found 04:42 pointto M5_0425 Tel="ACQUIRE" gs=0 PA=0 PAM-
CA=-10 IE= -40
04:43 pointto M5_0425 Tel="TRACK" gs=0 PA=0
04:44 takeseq 0.002 30
  • 044509 cube221
04:46 pointto M5_0444 Tel="TRACK" gs=0 PA=0 Elevation 26deg
04:47 takeseq 0.004 30
  • 044805 cube 222
04:48 pointto M5_0593 Tel="TRACK" gs=0 PA=0 Elevation 82deg
Star not found
04:54 pointto M5_0593 Tel="ACQUIRE" gs=0 PA=0 PAM-
CA=3.5 IE=-31
Focus has been changed too.
04:57 pointto M5_0593 Tel="TRACK" gs=0 PA=0
Coma strongly present.
04:58 pointto M5_0593 Tel="ACQUIRE" gs=0 PA=0 PAM-
Steve cleared the Active optics corrections and added some RX and RY
CA=11 IE=11
05:03 pointto M5_0593 Tel="TRACK" gs=0 PA=0
05 :05 takeseq 0.0004 30
  • 050507 cube 223
05:21 takepic 0.004
  • 052123

Collimation data

05:31 The integration time of the WFS has been changed to 15000ms.

05:34 pointto BS9150 Tel="ACQUIRE" gs=0 PA=0 PAM- Elevation 22deg
CA=9 IE=-15
  • Astigmatism ~2-3arcsec
05:41 pointto BS9150 Tel="ACTIVE" gs=0
05:46 takepic 1.0
  • 054644
05:47 pointto BS9153 Tel="ACQUIRE" gs=0 PA=0 PAM- Elevation 40deg
05:51 pointto BS9153 Tel="ACTIVE" gs=0
CA=3 IE=0
Steve cleared the C00s
05:54 takepic 2.0
  • 055433
05:55 pointto BS9158 Tel="ACQUIRE" gs=0 PA=0 PAM- Elevation 58deg
Steve cleared the C00s
05:59 pointto BS9158 Tel="ACTIVE" gs=0
05:59 takepic 2.0
  • 060000
06:00 pointto BS9166 Tel="ACQUIRE" gs=0 PA=0 PAM- Elevation 75deg
Steve cleared the C00s
CA=-3 IE=15
06:03 pointto BS9166 Tel="ACTIVE" gs=0
06:05 takepic 2.0
  • 060527
06:05 pointto BS9145 Tel="ACQUIRE" gs=0 PA=0 PAM- Elevation 32deg
Steve cleared the C00s
06:06 AdSec Arbitrator went offline and Armando has to force a restart by a command line.
  • Astigmatism ~4-5arcsec
06:31 pointto BS9145 Tel="ACTIVE" gs=0
06:34 takepic 2.0
  • 063529
06:36 pointto BS9162 Tel="ACQUIRE" gs=0 PA=0 PAM- Elevation 60deg
Steve cleared the C00s
The Vent Doors have been opened
06:41 pointto BS9162 Tel="ACTIVE" gs=0
3 06:43 takepic 2.0
  • 064332
06:43 pointto BS9158 Tel="ACQUIRE" gs=0 PA=0 PAM- Elevation 49deg
Steve cleared the C00s
  • Astigmatism ~2-3arcsec
CA=4 IE=6
06:46 pointto BS9158 Tel="ACTIVE" gs=0
06:51 takepic 2.0
  • 065100
06:51 pointto BS9169 Tel="ACQUIRE" gs=0 PA=0 PAM- Elevation 65deg
Steve cleared the C00s
CA=11 IE=4.5
06:54 pointto BS9169 Tel="ACTIVE" gs=0
06:58 takepic 2.0
* 065823
06:58 pointto BS9153 Tel="ACQUIRE" gs=0 PA=0 PAM- Elevation 25deg
Steve cleared the C00s
CA=-10 IE=-10
Focus has changed too
07:02 pointto BS9153 Tel="ACTIVE" gs=0
07:08 takepic 2.0
  • 070812

Transformation data

07:11 pointto StoneMagw Tel="ACTIVE" gs=0 Elevation 54.5deg
07:13 pointto StoneMagw Tel="ACQUIRE" gs=0 PA=0 PAM-
07:16 pointto StoneMagw Tel="ACTIVE" gs=0
07:18 pointto StoneMagw Tel="ACTIVE" gs=55
Seeing~0.6"
07:21 takepic 2.0
  • 072127
07:22 takepic 1.0
  • 072208
07:23 takepic 0.5
  • 072314
07:24 offste 30 0
07:25 takeseq 0.5 60
  • 072524 cube224
07:26 offset 0 0
07:27 cl< /home/LBTO/iraf/tran_StoneM_mix1.cl
  • first : 072836 cube225
  • last: : 081934 cube256
08:26 cl< /home/LBTO/iraf/tran_StoneM_mix2.cl Steve restarted Agw. (Issue Track 2804)
08:34 cl< /home/LBTO/iraf/tran_StoneM_mix2.cl
  • first : 083444 cube 257
  • last : 090737 cube276

Pointing model data

09:10 The integration time of the WFS has been changed to 5000ms.
09:11 gopoint exp-time=0.02s Nexp=500 delay=100s ploop+ log=gopointjuly062010
Steve cleared C00s and changed the focus between each star.
  • 091821 cube 277
  • 092528 cube 278
  • 093004 cube 279 : FAILED Elevation 26deg
  • 093422 cube 280
  • 093807 cube 281 : FAILED
  • 094258 cube 282
  • 094731 cube 283 : NOT VERY GOOD Elevation 33deg
  • 095151 cube 284
  • 095543 cube 285 : FAILED Elevation 20deg
  • 100007 cube 286
  • 100428 cube 287
  • 100928 cube 288 : FAILED Elevation 24deg
  • 101348 cube 289 : FAILED Elevation 39 deg
  • 101731 cube 290 : NOT VERY GOOD
  • 102122 cube 291 : FAILED
  • 102511 cube 292 : NOT VERY GOOD
  • 102943 cube293 : SATURATED
  • 103401 cube 294 : SATURATED
  • 103847 cube 295 : FAILED Elevation 16deg
  • 104248 cube 296
  • 104748 cube 297 : FAILED

Dice 5 data

10:50 pointto BS9188 Tel="ACQUIRE" gs=0 PA=0 PAM- Elevation 76deg
CA=2 IE= 20.5
10:57 pointto BS9188 Tel="ACTIVE" gs=29 PA=90
10:59 offset 30 0
11:00 takeseq 2.0 30
* 110004 cube 298
11:01 offset 0 0
11:01 dice5 5x5 RADEC 2.0 30 3 :
* first : 110307 cube 299
* last : 112645 cube 313

11:29 pointto BS9183 Tel="ACQUIRE" gs=0 PA=0 PAM- Elevation 48deg
CA=-12 IE=-7
11:36 pointto BS9183 Tel="ACTIVE" gs=29 PA=0
Very faint
11:38 pointto BS9185 Tel="ACTIVE" gs=23 PA=0
11:40 offset 30 0
11:40 takeseq 2.0 30
* 114040 cube 314
11:41 offset 0 0
Seeing ~0.73"
dice5 5x5 RADEC 2.0 30 3 :
  • first : 114401 cube 315
  • last : 115558 cube 323

Close
11:57 Close T-air=10.9C, Humidity=42%, Wind=1.2m/s
12:05 The Shell is rest.
12:06 Steve is moving the Hexapod of the secondary mirror at the positions given by Armando.
12:14 The M2 unit is power off and the GUIs have been closed.

-- PascaleHibon - 06 Jul 2010
Topic revision: r4 - 06 Jul 2010, PascaleHibon
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback